OpenStudy (albert0898):

SAT MATH 2 QUESTION WILL AWARD A MEDAL! If f(x)=|5-3x| then f(2)= A) f(-2) B) f(-1) C) f(1) D) f(4/3) E) f(7/3)

OpenStudy (vheah):

Subsititue the two in for x in f(x)=|3x-5| f(2)=|3(2)-5| = ?

OpenStudy (albert0898):

Answer is (D)... so your method is incorrect. I used your method at first and it was wrong. @vheah

OpenStudy (albert0898):

This is SAT Math 2...

OpenStudy (vheah):

I don't understand how it would be 4/3 if the meaning of f(2) means what would the function equal two if x is two. Where did you get the answer from?

OpenStudy (vheah):

what would the function equal t if x=2 * my bad.

OpenStudy (albert0898):

I figured out how to do it. Step 1. Plug 2 into the equation. You get |5-3(2)| = |-1| = 1 Step 2. This expression is now equivalent to 1. Solve for x. 1 = 5 - 3x -4 = -3x 4/3 = x @vheah

OpenStudy (vheah):

That makes perfect sense now. Good job!